### v2

# 3D Strain Field Reconstruction by Inversion of Dynamical Scattering

Abstract
Strain governs not only the mechanical response of materials but also their electronic, optical, and catalytic properties. For this reason, the measurement of the 3D strain field is crucial for a detailed understanding and for further developments of material properties through strain engineering. However, measuring strain variations along the electron beam direction has remained a major challenge for (scanning-) transmission electron microscopy (S/TEM). In this article, we present a method for 3D strain field determination using 4D-STEM. The method is based on the inversion of dynamical diffraction effects, which occur at strain field variations along the beam direction. We test the method against simulated data with a known ground truth and demonstrate its application to an experimental 4D-STEM dataset from an inclined pseudomorphically grown $\text{Al}_{0.47}\text{Ga}_{0.53}\text{N}$ layer.
The internal strain state governs the electronic (Thompson et al., 2006), optical (Jacobsen et al., 2006), and catalytic (Pingel et al., 2018) properties of materials, making its 3D measurement crucial for strain engineering. Applications range from virtually all modern CMOS-based field effect transistors (Cao et al., 2023), over strain rich electrocatalysts (Yang et al., 2022), to buried semiconductor quantum dots, where precisely tuned 3D strain profiles are essential to avoid fine structure splitting in entangled photon emission (Huber et al., 2018).
However, so far 3D strain fields have remained largely inaccessible to state-of-the-art strain measurement techniques in (scanning-) transmission electron microscopy (S/TEM). Several techniques exist like nano-beam electron diffraction (Usuda et al., 2005; Béché et al., 2009; Mahr et al., 2021), dark-field electron holography (Hÿtch et al., 2008; Béché et al., 2011), peak fitting (Bierwolf et al., 1993; Rosenauer et al., 1996), geometric phase analysis (Hÿtch et al., 2003), or other analyses of atomic resolution images (Niermann et al., 2011). Critically, all these techniques assume a strain field that is constant along the electron beam direction. Most of these techniques can only measure the strain components within the lateral plane, perpendicular to the electron beam, while the components in beam direction remain inaccessible for a measurement along a single direction. Furthermore, strain measurement techniques at atomic resolution often require thin specimen of a few ten nanometers at maximum, which alters the strain within the specimen due to surface relaxation effects (Jacob et al., 1998; Beyer et al., 2021).
For strain measurements the specimen must be oriented into strongly diffracting orientations, such as systematic row or zone-axis conditions. Under these orientations dynamical diffraction is dominant in samples thick enough to be representative of a bulk structure. For strain fields varying in beam direction dynamical diffraction results in a highly non-linear dependence of the measured signals on not only the strain inhomogeneity and its depth, but also on parameters like specimen thickness and the beam tilt (Lubk et al., 2014; Javon et al., 2014; Meißner et al., 2019), further complicating strain analysis with most techniques. In contrast to electrons, X-rays do not exhibit strong dynamical diffraction effects, which enables advanced computational techniques like Bragg coherent diffraction imaging to be used for 3D strain measurements (Pfeifer et al., 2006; Cherukara et al., 2018).
In this article, we introduce an inversion method that determines the 3D strain field from 4D-STEM data. This inversion method measures the strain not only in lateral directions, but also in the electron beam direction. The basic idea of this method is to utilize the rich information encoded in dynamical diffraction patterns to solve the inverse problem for the 3D strain field. As input we acquire scanning convergent beam electron diffraction (SCBED) datasets (Niermann et al., 2025), where the electron probe is scanned over the specimen and convergent beam electron diffraction (CBED) patterns are acquired at each scan position. The beam convergence angle is kept smaller than the Bragg angle to avoid disk overlap within the CBED patterns. SCBED is a special 4D-STEM technique that is especially useful for measuring dynamical diffraction effects from 3D strain fields (Niermann et al., 2024; Otto et al., 2025). The results obtained by the inversion method are spatial mappings of the strain field not only in the lateral dimension but also along the electron beam direction. The method itself is widely applicable, since these SCBED datasets can be recorded with any 4D-STEM capable microscope and the numerical reconstruction itself can be performed on modern desktop computers.
Below, we introduce the reconstruction method in detail, test it against simulated data with known ground truth, and finally demonstrate the reconstruction of the strain field variations along the beam direction on an experimental 4D-STEM dataset from an inclined pseudomorphically grown $\text{Al}_{0.47}\text{Ga}_{0.53}\text{N}$ layer.
To simplify the description, we consider lateral effects only along the lateral $x$-direction within this article. The $z$-direction corresponds to the electron beam direction. Furthermore, we only consider zeroth order Laue Zone scattering within systematic row diffraction conditions. Under systematic row conditions, all excited crystal reflections are oriented along a single direction (De Graef, 2003), which makes the diffraction pattern essentially one-dimensional. Here, the systematic row is parallel to the $x$-direction. This geometry suffices for the presented cases. However, this is no general restriction of this method and its extension to the $y$-direction, different diffraction geometries, and a more general set of reflections is straightforward.
For the systematic row geometry, all information of the CBED pattern can be obtained from an one-dimensional profile (with $q$ being the coordinate along this profile) through the center of the CBED disks along the systematic row. We can display the intensities of these diffraction profiles $I(x,q)$ for each spatial $x$-coordinate in a two dimensional $qx$-plot (Niermann et al., 2025). Each row of this $qx$-plot corresponds to the diffraction profile obtained at the corresponding position $x$. Experimental 4D-STEM datasets can be reduced to such $qx$-plots $I^{(\mathrm{exp})}(x,q)$ by calculating profiles along the corresponding directions. Examples for such $qx$-plots can be seen below, for instance in Fig. 4(b). Fig. 4(c) shows exemplarily the CBED pattern, from which the row at $x=9$ nm in the $qx$-plots of Fig. 4(b) is derived.
The deformation of crystals can be described either by its displacement field $\vec{u}(\vec{r})$ or by the strain tensor. The strain tensor is defined as the symmetric part of the Jacobian of the displacement field, while its antisymmetric part describes an infinitesimal rotation of the crystal. Nevertheless, we will use the term “strain” for the $z$-derivatives of the $x$-component of the displacement field in the following:
$\epsilon(x,z)=\frac{\partial}{\partial z}u_{x}(x,z).$
We reconstruct the strain field $\epsilon(x,z)$ numerically as follows: We choose a regular rectilinear grid with step sizes $\delta x$ and $\delta z$ for discretization:
$\epsilon_{vw}=\epsilon(v\delta x,w\delta z)$
with integer indices $v$ and $w$. Evaluations of the strain field $\epsilon(x,z)$ at positions in between the grid points are performed by bilinear interpolation of the values of the surrounding grid points.
For a given specimen thickness $t$, and a given strain field $\epsilon(x,z)$ the expected complex-valued beam amplitudes $\phi_{g}$ for the reflections $g$ can be calculated within the validity of the column approximation for each $x$-position and each lateral component $k_{\perp}$ of the incident beam vector by numerical propagation of the Darwin-Howie-Whelan (DHW) equations (De Graef, 2003; Williams and Carter, 2009):
$\frac{\partial}{\partial z}\phi_{g}(z;x,k_{\perp})=\mathrm{i}2\pi\left[s_{g}+g\epsilon(x,z)\right]\phi_{g}(z;x,k_{\perp})+\\ \mathrm{i}\frac{\pi}{k_{0}}\sum_{g^{\prime}}U_{g-g^{\prime}}\phi_{g^{\prime}}(z;x,k_{\perp}),$ (1)
with the excitation error
$s_{g}=-\frac{g\cdot(g+2k_{\perp})}{2k_{0}},$
the structure factors $U_{g}$ of the investigated material, which we calculate using absorptive atomic form factors(Weickenmeier and Kohl, 1991), and the vacuum wave number $k_{0}.$ In this work, a 4th order Runge-Kutta scheme with a step size of $0.2$ nm was used for this propagation and no material variations were considered within the reconstructions.
The intensity of a single reflected beam in the exit plane $z=t$ is given by the modulus square of the beam amplitude
$I_{g}(x,k_{\perp})=\left\|\phi_{g}(z\equiv t;x,k_{\perp})\right\|^{2}.$ (2)
and the intensities $I(x,q)$ observed in the corresponding $qx$-plot are given by (for a residual beam tilt of $\Theta)$:
$I(x,q)=\sum_{g}I_{g}(x,q-g+k_{0}\Theta)w(q-g).$ (3)
The weighting function $w(q)$ is used to mimic the effects of the condenser aperture with a semiconvergence angle of $\Theta_{\text{max}}$, but replaces the hard edges of the aperture by the smooth sigmoid function $\textrm{sig}(q)$ to avoid artifacts:
$w(q)=\mathrm{sig}\left(\beta(k_{0}\Theta_{\mathrm{max}}-q)\right)\mathrm{sig}\left(\beta(k_{0}\Theta_{\mathrm{max}}+q)\right)\\ \text{ \,\,with\,\, }\mathrm{sig}(q)=\frac{1}{1+\exp(-q)},$
where $\beta$ is a softening parameter for the edges.
Having described the physical forward model, we now turn to the inverse problem. The goal of our inversion method is the reconstruction of the strain field (i.e. the values for $\epsilon_{vw}$) from $qx$-plots $I^{\mathrm{(exp)}}(x,q)$ of experimentally measured data. This can be achieved by numerical minimization of the loss function $L[\epsilon_{vw}]$ with respect to $\epsilon_{vw}$:
$L[\epsilon_{vw}]=R[\epsilon_{vw}]+\lambda S[\epsilon_{vw}].$ (4)
The residual norm
$R[\epsilon_{vw}]=\sum_{gxq}\left[w(q-g)\left(NI_{\vec{g}}(x,q-g+k_{0}\Theta)-I^{\text{(exp)}}(x,q)\right)\right]^{2}$ (5)
drives the solution towards a strain field that accurately reproduces the experimental data. The regularization term
$S[\epsilon_{vw}]=\sum_{vw}\epsilon_{vw}^{2},$
weighted by $\lambda$, penalizes non-physical, noisy solutions and promotes a physically plausible, smooth strain field. The parameter $N$ adjusts the simulated illumination strength to best match the experimental data by minimizing $R[\epsilon_{vw}]$. The $\mathcal{L}^{2}$-norm (Tikhonov regularization) is used as regularization term, motivated by the expectation that the deformation energy will scale with the square of the strain. The regularization strength $\lambda$ is chosen using the L-curve criterion (Hansen, 2001). In all cases presented in this article, the residual norm (5) is calculated only over the data points shown in the $qx$-plots.
The strain field was determined by numerically minimizing the loss function (4) using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm with automatically calculated gradients. The minimization was initialized with a zero strain field $\epsilon_{vw}\equiv 0$.
To first demonstrate the fundamental feasibility of the approach, we tested it on a simulated dataset with a known ground truth strain field, namely the 3D strain field of an $\text{In}_{0.4}\textrm{Ga}_{0.6}\textrm{As}$ spherical precipitate, here modeled as the continuum strain model of a hard sphere (Williams and Carter, 2009). This test case is an example for specimens with intrinsic 3D strain fields as not only found for precipitates but also in buried quantum dots or other inclusions of different materials. The precipitate has a radius of 6 nm and is located 25 nm behind the entrance surface in a 95 nm thick GaAs specimen. The specimen is oriented in $\{220\}$-systematic row conditions ($g$-direction). The $x$-direction is parallel to the direction of the systematic row. The assumed acceleration voltage is 300 kV. Fig. 1(a) shows the theoretical strain field $\epsilon_{\text{Theo}}(x,z)$ (relative to the embedding GaAs). It shows the four characteristic lobes of the shear strain, which are typical for spherical precipitates (Williams and Carter, 2009).
Figure 1: (a) Theoretical strain field $\epsilon_{\text{Theo}}(x,z)$ of a spherical precipitate located at a depth of 25 nm. (b) Simulated qxqx-plot for the strain field $\epsilon_{\text{Theo}}(x,z)$ in (a). (c) Strain field $\epsilon(x,z)$ reconstructed by the inversion method from the simulated qxqx-plot in (b). (d) Simulated qxqx-plot for the reconstructed strain field $\epsilon(x,z)$ in (c).
The $qx$-plot shown in Fig. 1(b) was simulated by numerical propagation of the DHW equations for the theoretical strain field of Fig. 1(a) for a scan across the center of the precipitate. It shows the intensity variations originating from dynamical diffraction at the strain field of the precipitate. A semi-convergence angle of 4.3 mrad was used.
The inversion method is applied to reconstruct the strain field $\epsilon(x,z)$ from this $qx$-plot. The results are shown in Fig. 1(c). The reconstructed strain shows a lower resolution, especially in the $z$-direction, as can be seen in the slightly smeared out strain field in beam direction. Nevertheless, the inversion method has determined the strain field and the depth of the precipitate solely from the information encoded within the dynamical diffraction patterns in Fig. 1(b). Fig. 1(d) shows the $qx$-plot simulated for the reconstructed strain, which exhibits a very good match to the $qx$-plot in Fig. 1(b).
The position of the precipitate as well as the four characteristic lobes are clearly reconstructed. While the position could also be determined from other methods (e.g. by stereoscopy), the reconstruction of these four lobes provides insights into the shape of the precipitate, which were previously not accessible. The actual size of a precipitate is usually not easily determinable from S/TEM data, as the dominating contrasts originate from dynamical diffraction of the 3D strain field and not from the precipitate itself. As strain effects are typically strongest at the interfaces, the 3D position of the strain features already give a reliable idea of the precipitate’s extents and shape. Furthermore, stronger localized strain features typically originate from sharp points in the shape, while extended features like the observed 4 lobes are typical for precipitates with smooth shapes.
A key aspect of any dynamical diffraction-based method is understanding its intrinsic symmetries. To investigate this, we applied the method to a simulated inclined layer, a geometry known to produce symmetric diffraction effects. The results are shown in Fig. 2. The assumed specimen contains an inclined $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer of 5 nm thickness embedded within a 115 nm thick GaN matrix. This application is an example for strain fields in specimen with extended structures as found for instance in semiconductor heterostructures, but also is prototypical for extended defects like stacking faults. The $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer is located within the basal plane of the Wurtzite structure. The specimen is oriented in $\{01\overline{1}2\}$-systematic row conditions. The $x$-direction is parallel to the systematic row direction. The crystal’s $c$-axis lays within the $xz$-plane with an inclination of 43° with respect to the $x$-direction. The assumed acceleration voltage is 200 kV. Fig. 2(a) shows the theoretical strain field $\epsilon_{\text{Theo}}(x,z)$ of the inclined layer (relative to the GaN host material).
The $qx$-plot shown in Fig. 2(b) was simulated by numerical propagation of the DHW equations for the theoretical strain field of Fig. 2(a). A semiconvergence angle of 6.3 mrad was used. The $qx$-plot in Fig. 2(b) shows a mirror symmetry in $x$-direction with respect to $x\approx 67$ nm, which is the position where the inclined layer crosses the specimen’s midplane.
Figure 2: (a) Theoretical strain field $\epsilon_{\text{Theo}}(x,z)$ of an inclined $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer. (b) Simulated qxqx-plot for the strain field $\epsilon_{\text{Theo}}(x,z)$ in (a). (c) Strain field $\epsilon(x,z)$ reconstructed by the inversion method from the simulated qxqx-plot in (b). (d) Simulated qxqx-plot for the reconstructed strain field $\epsilon(x,z)$ in (c).
The inversion method is applied to reconstruct the strain field $\epsilon(x,z)$ from this $qx$-plot. The resulting reconstructed strain field $\epsilon(x,z)$ is shown in Fig. 2(c). The reconstructed strain field $\epsilon(x,z)$ in Fig. 2(c) and the theoretically strain field $\epsilon_{\text{Theo}}(x,z)$ in Fig. 2(a) differ significantly, even though the corresponding simulated $qx$-plots Fig. 2(b) and Fig. 2(d) are nearly identical.
This is a direct consequence of a well-known symmetry of dynamical diffraction (Koprucki et al., 2022; Pogany and Turner, 1968; Howie and Whelan, 1961): a mirroring of the strain field $\epsilon(x,z)$ with respect to the specimens midplane $t/2$ results in the same $qx$-plots
$I[\epsilon(x,z)]=I[\epsilon(x,t-z)].$
This symmetry was also already observed in experimental $qx$-plots of inclined layers (Niermann et al., 2025) and dislocations (Niermann et al., 2024).
This symmetry leads to an ambiguity in the reconstructed strain profiles. In principle, this ambiguity might lead to an arbitrary flipping of the reconstructed $\epsilon(x,z)$ with respect to $t/2$ between neighboring $x$ points. However, the used strain discretization with a coarse $x$-spacing stabilizes the reconstructions with a continuous $\epsilon(x,z)$ in $x$-direction. Both effects combined lead to the four compatible continuous reconstructions of the strain fields $\epsilon(x,z)$ that are shown in Fig. 3(a). All four strain fields will result in nearly the same intensities within the $qx$-plots, thus are mathematically correct solutions to the inversion problem. The reconstruction method converges on one of the possible solutions in Fig. 3(a). Which of the four cases eventually is reconstructed will depend on the starting conditions (and in experimental data also on noise and other experimental influences).
However, this mathematical ambiguity is not a practical impasse. Only one of these solutions is typically physically plausible, therefore a unique solution can almost always be identified through minimal physical or crystallographic considerations: Since the $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer has smaller lattice constants than the embedding GaN, the displacement field component in the crystalline $[0001]$-direction (the Wurtzite’s $c$-axis) must decrease in this direction as the elongation along the $c$-axis is negative. However, two possibilities for the orientation of the basal plane exist, which are compatible with a $\{01\overline{1}2\}$-normal in $x$-direction. These are shown in Fig. 3(b) and 3(d). Their resulting derivatives of the displacement field along the $z$-direction are shown in Fig. 3(c) and 3(e). However, only the one in Fig. 3(e) will exhibit a negative derivative within the layer, making this the only physically possible solution.
Figure 3: (a) Sketches of strain fields $\epsilon(x,z)$, which will result in the same qxqx-plots. These four strain fields are obtained by flipping the left half, the right half, or both of them at the specimen midplane at $z=t/2$. (b) Displacement field $u_{x}$ with $[0001]$-direction inclined towards the top surface. (c) Derivative $\partial u_{x}/\partial z$ of the displacement field in (b). (d) Displacement field $u_{x}$ with $[0001]$-direction inclined towards the bottom surface. (e) Derivative $\partial u_{x}/\partial z$ of the displacement field in (d).
Having tested the method using simulations, we now demonstrate its application to a real-world experimental dataset as shown in Fig. 4. The experimental data was obtained from a zero-loss filtered 4D-STEM measurement of an inclined $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer of $5.0\pm 0.5$ nm thickness embedded within a $52.5\pm 2.5$ nm thick GaN lamella. This specimen was chosen, as it allows for a controlled strain field varying in the depth of the specimen. The layer was grown on the basal plane of the Wurtzite structure and was previously characterized in on-edge orientation. The inclination was realized by preparation of a lamella under an angle of 43° with respect to the growth direction by means of focused ion beam milling (Meißner et al., 2019) (see supplementary material for details).
The specimen is oriented in $\{01\overline{1}2\}$-systematic row conditions and the $x$-direction is along the $\{01\overline{1}2\}$-normal. The crystal’s $c$-axis lays within the $xz$-plane with an inclination of 43° with respect to the $x$-direction. The expected strain field for this specimen under the assumption of a pseudomorphic growth(Pohl, 2013) and without the consideration of surface relaxations is shown in Fig. 4(a). The measurement was performed on TEAM I, a double-aberration-corrected FEI Titan 80-300 microscope, operating at 300 keV. The 4D-STEM data was acquired using the GIF camera with a binning of 4 (for further acquisition details see the supplementary material).
Fig. 4(b) shows the $qx$-plot obtained from this measurement. The experimental semiconvergence angle was 5.0 mrad, which corresponds to a probe size of roughly 0.4 nm. Clearly visible is the change within the $qx$-plots for $x$-coordinates between -25 nm and 25 nm, which is caused from the scattering at the inclined layer. Furthermore, a slight shifting of the reflections itself towards negative $q$-coordinates with increasing $x$-coordinates can be seen within the measurement. This is a measurement artifact caused by a misalignment of the beam shift-tilt decoupling within the microscope. Furthermore, a slight shift of the intensity patterns within the reflections to negative $q$ coordinates with increasing $x$-coordinates can be noticed. This is caused by a slight bending of the specimen (Niermann et al., 2025). Both artifacts have been corrected by incorporating an additional shear transformation within the $qx$-plane into the reconstruction. The lamella thickness $t=52.5\text{\penalty 10000\ nm}$ and the residual beam tilt were evaluated in advance by minimizing the loss function with respect to these parameters in the unstrained areas.
Figure 4: (a) Expected strain field for a pseudomorphically grown 5.0 nm thick $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer. (b) Experimentally obtained qxqx-plot from the inclined layer. (c) Exemplary CBED pattern, with indicated profile direction, which corresponds to the row at $x=9$ nm indicated in (b). (d) Strain field $\epsilon(x,z)$ reconstructed from data in (b). (e) Calculated qxqx-plot for reconstructed strain field in (d).
The inversion method is applied to reconstruct the strain field $\epsilon(x,z)$ from this $qx$-plot. The result is shown in Fig. 4(d). Within Fig. 4(d) the inclined layer is the prevailing feature and it well reproduces to the strain field expected for the layer, which is shown in Fig. 4(a). The reconstructed strain field in Fig. 4(d) represents one of the mathematically valid solutions consistent with the symmetries discussed above. We identify this as the physically correct solution based on the known expected compressive strain in the AlGaN layer. The excellent agreement between the $qx$-plot Fig. 4(e) simulated for the reconstructed strain field and the experimental data in Fig. 4(b) – despite the presence of shot noise and minor experimental artifacts – demonstrates the practical robustness of the method.
We demonstrated the reconstruction of 3D strain fields by the presented reconstruction method. However, it is important to understand its current limitations and potential sources of error. We identify three main categories of uncertainty: (i) fundamental ambiguities arising from physical symmetries, (ii) systematic errors from inaccuracies in experimental parameters such as specimen thickness and residual tilt and (iii) statistical uncertainty from experimental noise.
The primary limitation is the inherent ambiguity due to the intrinsic symmetry of dynamical diffraction. A key direction for future work will be the implementation of strategies for breaking this symmetry, removing the need for post-reconstruction analysis and paving the way for a fully automated 3D strain mapping workflow. One option to overcome this ambiguity is the use of additional regularizations, which favors the physically correct solution. Another option is to evaluate the beam phases, which, in contrast to the intensities, are not susceptible to the symmetry. This can be achieved either by utilizing an overlap between neighboring CBED disks like also used in ptychography or by the acquisition of holographic tilt-series instead of STEM-data. A third option is the incorporation of additional measurements taken with rotated specimen into the inversion method. Incorporating multiple projection directions also will enable the 3D mapping of all components of the strain tensor.
We mitigated the influence of systematic errors, by determining these parameters from strain-free reference regions of the sample prior to the main reconstruction. The camera length (or more specifically the reciprocal space sampling of the CBED patterns) as well as the effects of misaligned decoupling of beam tilt and shift, can be calibrated from the positions of the reflections in these reference regions, as it is also standard in state-of-the-art STEM based measurement methods. Residual tilt and thickness can be estimated by comparing the CBED intensity patterns in these areas with simulations. The influences of these systematic effects can also be accounted for within the reconstruction: we also added the possibility to our method to optimize for the specimen thickness and reciprocal space sampling in addition to the strain field. However, we noticed no significant differences in the presented reconstructions. Residual tilt of the specimen will occur as a constant background in the reconstructed strain, as infinitesimal rotations of the specimen contribute to the derivative of the displacement field. Eventually this is more a question of relative to which reference the strain is measured. In principle defocus can be accounted for as an additional shear of the dataset within the $qx$-plane (Niermann et al., 2025).
As demonstrated in the reconstruction from experimental data, shot noise is no limitation in the present example, but obviously will influence the quality of the reconstructions under lower dose conditions.
Future work will focus on quantifying the method’s precision by systematically investigating the impact of noise and other experimental influences on the reconstructed strain. The demonstrated performance of the method under standard acquisition conditions suggest that high-fidelity reconstructions are achievable.
In conclusion, we have developed and demonstrated an inversion method that reconstructs 3D strain fields from 4D-STEM data, providing access to strain variations along the electron beam direction - a capability unavailable to existing S/TEM techniques. Our technique is not confined to thin specimens; it is specifically designed for thicker samples (>50 nm), where it can measure strain variations deep within the material, free from the surface relaxation effects that challenge conventional strain measurement techniques. Furthermore, the method does not require a-priori models of the strain distribution.
Our method provides a direct solution to the long-standing challenge of quantifying 3D strain fields in technologically relevant systems, such as catalysts and (opto-) electronic semiconductor devices. By inverting dynamical diffraction effects, we can now, for instance, resolve the size, shape, and strain state of buried features like quantum dots and precipitates in 3D – information previously obscured by the effects of dynamical diffraction. The determination of 3D strain fields is immediately applicable to optimizing the optoelectronic properties of semiconductor heterostructures, understanding failure mechanisms in structural alloys, and engineering the catalytic activity of core/shell nanoparticles. By providing a quantitative, depth-resolved picture of strain, this technique enables a more precise correlation between nanoscale structure and device performance, offering a powerful tool for materials design and process control.
Supplementary Material
See supplementary material for a more detailed description of the investigated specimen and its preparation, and for further parameters of the experiment and the reconstruction.
Acknowledgements.
L. N. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project 492463633. Work at the Molecular Foundry was supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. We thank the Center for Electron Microscopy (ZELMI), Technische Universität Berlin, and the Structure Research and Electron Microscopy group, Humboldt-Universit"at zu Berlin, in the context of the Alliance Center Electron Microscopy (ACEM) for support. The ACEM is funded under the Excellence Strategy of the Federal Government and the Länder by the Berlin University Alliance (BUA).
Data Availibility Statement
The data that support the findings of this study are openly available in the Zenodo repository at https://doi.org/10.5281/zenodo.15309273(Niermann, 2025)
Author Declarations
The authors have no conflicts to disclose.
Author Contributions
Laura Niermann: Conceptualization, Funding Acquisition, Investigation, Methodology, Project Administration, Resources, Software, Supervision, Writing – original draft, Writing – review & editing. Tore Niermann: Investigation, Visualization, Data Curation, Software, Validation, Writing – review & editing. Cheyun Song: Investigation, Writing – review & editing. Colin Ophus: Resources, Validation, Writing – review & editing.
Supplemental Materials: 3D Strain Field Reconstruction by Inversion of Dynamical Scattering
S1 Details on experiment
The STEM lamella, from which the experimental data was obtained, was prepared by focused ion beam (FIB) milling from a semiconductor heterostructure. The heterostructure consists out of a 5 nm thick $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer embedded within GaN. The layer structure was grown on (0001) planes using metal-organic chemical vapor deposition. Thickness and concentration of the layer were verified by means of EDX-STEM on a separate cross-section STEM-lamella. For the inclined layer structure investigated in this work the lamella was prepared under an angle 43° with respect to the wafer’s (0001) surface as sketched in Fig. S1(a).
Figure S1: FIB preparation of the inclined layer structure: (a) The FIB lamella preparation was performed under an angle of 43° with respect to the wafer surface and the $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer. (b) The resulting STEM lamella contains the inclined layer.
After FIB preparation the lamella was transferred into the STEM, where the lamella is observed with the lamella surfaces nearly perpendicular to the beam. The specimen is tilted within the STEM a few degrees around the $(01\overline{1}2)$-normal to achieve $(01\overline{1}2)$-systematic row conditions. In this orientation the specimen eventually contains an inclined $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer with a well known geometry as shown in Fig. S1(b).
The 4D-STEM data was obtained with a scan-stepping of 0.31 nm. The sampling of the diffraction pattern was 0.11 mrad/px. For the $qx$-plots, at first profiles in spatial dimensions were obtained across the inclined layer. Fig. S2 shows an annular dark field overview over the inclined layer. At the top of the image the protective Pt-layer from FIB preparation is visible. Within the region of the projected inclined layer the image contrast is slightly decreased. For better visibility the upper and lower projected edge are indicated by a dashed line in Fig. S2. The position of the aforementioned profile is indicated by the arrow. The individual diffraction patterns of the 4D-STEM dataset were averaged over a window of 16 px width in the perpendicular direction (parallel to the inclined layer) to increase the signal-to-noise ratio. This results in a 3D datasets (an averaged 2D diffraction pattern for each spatial position $x$).
Figure S2: Annular dark field overview image over the inclined $\mathrm{Al}_{0.47}\mathrm{Ga}_{0.53}\textrm{N}$-layer. The dashed lines mark the projected edges of the inclined layer. The arrow indicates the region over which the profile within the spatial dimensions was obtained. The position $x$ within the $qx$-plots corresponds to a point on this profile. In the top part of the image the protective Pt layer can be seen.
These diffraction patterns are shown for exemplary $x$-positions in Fig. S3. The dataset was further reduced by obtaining profiles within the diffraction patterns along the systematic row in the central part of the CBED disks. The arrows in Fig. S3(b) mark the position of this profile. These profiles were averaged over 9 px in the direction perpendicular to the arrow. This averaging was required to average out the influence of the HOLZ-lines noticeable within the $(0000)$-diffraction disk.
Figure S3: Exemplary CBED patterns for several $x$-positions. (a) The qxqx-plots with marked x-positions, for which the CBED patterns are shown. (b) CBED patterns for the indicated $x$-positions. The arrow indicates, where the profile was obtained. The intensities within the $qx$-plot and the individual diffraction patterns are displayed on a common color scale, which is given in (a)
After obtaining this second profile, a 2D dataset is left, which is displayed within the $qx$-plot as shown in Fig. S3(b) (this is the same $qx$-plot as shown in Fig. 4(b) of the main text). For each position $x$ within this plot the profile of the corresponding diffraction pattern is shown along the $q$-position.
S2 Experimental and reconstruction parameters
Table S1 gives the detailed parameters used in the simulations, experiments and reconstructions.
Simulated Simulated Experimental
Parameter precipitate inclined layers inclined layer
Semiconvergence angle (mrad) 4.3 6.3 5.0
Residual beam tilt (mrad) 0.0 0.0 0.5
$x$-sampling qxqx-plot (nm) 2.0 2.0 0.31 (see text)
$q$-sampling qxqx-plot (1/nm) 0.1 0.1 0.057
Regularization strength $\lambda$ 125 $2\times 10^{4}$ $3\times 10^{9}$ and
Max. half-angle $\Theta_{\text{max}}$ (mrad) 4.3 6.3 4.4 (see text)
Smoothing parameter $\beta$ (nm) 20 20 10
Grid spacing $\delta x$ (nm) 2.56 6.74 7.14
Grid spacing $\delta z$ (nm) 2.1 10.45 6.87
Table S1: Parameters in experiments and reconstructions
The qxqx-plots within the experimental inclined layer shows some diffraction effects at the condenser aperture edges. For avoidance of reconstructions artifacts that originate from these effects, a parameter of $\Theta_{\mathrm{max}}=4.4\thinspace\text{mrad}$ was used, which is slightly smaller than the experimental semiconvergence angle. During the reconstruction of the strain field a coarser $x$-spacing of only $2$ nm was used within the calculated qxqx-plots for efficiency. For the evaluation of the residual norm these coarser points within $I(x,q)$ were linearly interpolated in the $x$-direction. The $q$-spacing was identical to the spacing in the experiment.